Ahmed has budgeted next year’s sales at 8,000 units. (2 marks )

Compute Ahmed's degree of operating leverage. If P = 1,000, V = 400, F = 850,000.
Answer #1

Ans-DOL=Contribution Margin/ Operating Income

=$8, 850,000/$8,000,000

=1.11 times

Contribution margin = Profit + fixed cost

=1,000*8,000=8,000,000+850,000

=8,850,000

Operating income= 1,000*8,000=8,000,000